OpenStudy (anonymous):

The speed, s, of an ant (in centimeters per second) is related to the temperature, t (in degrees Celsius), by the formula s=0.167t-0.67 If an ant is moving at 4 centimeters per second, what is the temperature? Round your answer to the nearest degree.

OpenStudy (mathstudent55):

Plug in 4 for s and solve for t.

OpenStudy (anonymous):

s=4 find t

OpenStudy (anonymous):

-0.2

OpenStudy (anonymous):

@Jonask

OpenStudy (anonymous):

4=0.167t-0.67 4.67=0.167t t=4.67/0.167=27.96